A projectile is launched from and returns to ground level. Air resistance is absent. The horizontal range of the projectile is measured to be R = 175 m, and the horizontal component of the launch velocity is Vox = 25 m/s. Find the vertical component Voy of the launch velocity.
1
Expert's answer
2019-09-20T09:16:58-0400
"R=\frac{v^2\sin{2\theta}}{g}""V_{0x}=vcos{\theta}=25"

V_{0y}=vcos{\theta}

"V_{0y}=vsin{\theta}"

"V_y=v\cos\theta=17\frac{m}{s}"